Give an ordered pair (x,y) of numbers that satisfy the equation
4x+y=3

Answer by Earlsdon(6294)

Rewrite the equation in the slope-intercept form: y = mx+b
y = -4x+3 Now substitute any value you like for x and you will get the corresponding value of y. Let x = 5
y = -4(5)+3
y = -20+3
y = -17
So one ordered pair would be (5, -17)
As you can see, there are an infinite ordered pairs that would satisfy this equation, you choose the x-value and there will be a corresponding y-value..